# Retrieval-Augmented Layout Transformer for Content-Aware Layout Generation

### v3

@@ 4.3 Baseline Methods
${}^{\dagger}$ [19] is a discrete state-space diffusion model that can handle many constrained generation tasks. Since the model is originally designed for content-agnostic layout generation, we extend the model to accept an input image.
@@ 4.4 Implementation Details
${}^{\dagger}$ 0.150 0.0192 0.41 0.190 27.09
${}^{\dagger}$ 0.123 0.0144 0.51 0.091 10.03
Table 3: Retrieval augmentation for CGL-GAN and ${}^{\dagger}$ on the PKU test split.
@@ 4.5 Unconstrained Generation
While our RALF is an autoregressive generator, we show that retrieval augmentation also benefits other generative models for content-aware layout generation. Here, we adapt CGL-GAN and ${}^{\dagger}$ with retrieval augmentation and evaluate the performance. Table 3 summarizes the results. CGL-GAN and ${}^{\dagger}$ combined with our retrieval augmentation consistently improve many evaluation metrics. We provide additional results in the appendix.
@@ 4.6 Constrained Generation
Category normal-$\rightarrow$ Size + Position (C normal-$\rightarrow$ S + P) takes in element types and generates the sizes and positions for each element.
Category + Size normal-$\rightarrow$ Position (C + S normal-$\rightarrow$ P) generates element positions based on given element categories and sizes.
@@
Figure 7 summarizes constrained generation results. The results indicate that RALF is effective even for constrained generation tasks. For tasks such as C + S $\rightarrow$ P and Refinement, RALF shows notable improvement in the $\mathrm{FID}$ metric. This suggests that referencing authentic examples to understand element relationships enhances position prediction accuracy. Overall, the results highlight RALF’s capability to significantly augment the generative performance over the baseline approach.
@@ D Additional Results
CGL-GAN ${}^{\dagger}$ Autoreg Baseline RALF
@@
[…]
@@
Table F shows the results of retrieval augmentation for CGL-GAN and ${}^{\dagger}$. Even for constrained generation tasks, retrieval augmentation achieves a better quality of generation for other generators on almost all metrics.
@@
${}^{\dagger}$ 0.115 0.0121 0.57 0.008 2.86 0.127 0.0188 0.75 0.018 1.98
${}^{\dagger}$ 0.115 0.0121 0.57 0.007 2.91 0.126 0.0186 0.76 0.019 1.79
Table F: Retrieval augmentation for CGL-GAN and ${}^{\dagger}$ on the PKU and CGL test split for unconstrained and constrained generation.
@@
Autoreg ${}^{\diamondsuit}$ 256 6.59M 0.134 0.0165 0.44 0.018 13.51 0.125 0.0190 0.92 0.011 2.90
${}^{\diamondsuit}$ 0.119 0.0129 0.92 0.008 3.45 0.125 0.0180 0.98 0.004 1.31

### v4

@@ 4.3 Baseline Methods
LayoutDM [19] is a discrete state-space diffusion model that can handle many constrained generation tasks. Since the model is originally designed for content-agnostic layout generation, we extend the model to accept an input image.
@@ 4.4 Implementation Details
LayoutDM 0.150 0.0192 0.41 0.190 27.09
LayoutDM 0.123 0.0144 0.51 0.091 10.03
Table 3: Retrieval augmentation for CGL-GAN and LayoutDM on the PKU test split.
@@ 4.5 Unconstrained Generation
While our RALF is an autoregressive generator, we show that retrieval augmentation also benefits other generative models for content-aware layout generation. Here, we adapt CGL-GAN and LayoutDM with retrieval augmentation and evaluate the performance. Table 3 summarizes the results. CGL-GAN and LayoutDM combined with our retrieval augmentation consistently improve many evaluation metrics. We provide additional results in the appendix.
@@ 4.6 Constrained Generation
Category $\rightarrow$ Size + Position (C $\rightarrow$ S + P) takes in element types and generates the sizes and positions for each element.
Category + Size $\rightarrow$ Position (C + S $\rightarrow$ P) generates element positions based on given element categories and sizes.
@@
Table 5 summarizes constrained generation results. The results indicate that RALF is effective even for constrained generation tasks. For tasks such as C + S $\rightarrow$ P and Refinement, RALF shows notable improvement in the $\mathrm{FID}$ metric. This suggests that referencing authentic examples to understand element relationships enhances position prediction accuracy. Overall, the results highlight RALF’s capability to significantly augment the generative performance over the baseline approach.
@@ D Additional Results
CGL-GAN LayoutDM Autoreg Baseline RALF
@@
Tables D and E present the quantitative result on the annotated test split without user constraints on the PKU and CGL datasets, respectively. RALF notably improves Density and Coverage metrics, indicating that RALF can generate better layouts in terms of both fidelity and diversity. RALF does not achieve the best score regarding Rshm and Alignment. However, these metrics may not be very reliable since the best scores for these metrics largely deviate from the scores for Real-Data, unlike other metrics.
@@
Table F shows the results of retrieval augmentation for CGL-GAN and LayoutDM. Even for constrained generation tasks, retrieval augmentation achieves a better quality of generation for other generators on almost all metrics.
@@
LayoutDM 0.115 0.0121 0.57 0.008 2.86 0.127 0.0188 0.75 0.018 1.98
LayoutDM 0.115 0.0121 0.57 0.007 2.91 0.126 0.0186 0.76 0.019 1.79
Table F: Retrieval augmentation for CGL-GAN and LayoutDM on the PKU and CGL test split for unconstrained and constrained generation.
@@
Autoreg Baseline 256 6.59M 0.134 0.0165 0.44 0.018 13.51 0.125 0.0190 0.92 0.011 2.90
RALF 0.119 0.0129 0.92 0.008 3.45 0.125 0.0180 0.98 0.004 1.31